The directions say "Use product property to simplify the expression".
1.) Radical 240

Answer by checkley75(3666)

Product property??????
sqrt240=sqrt2*2*2*2*15=sqrt4*4*15=4sqrt15.
------------------------------------------------------------
I hope this is what you need!